OpenStudy (anonymous):

Find the point that is one-fourth of the way from P(1,8) to Q(-7,8)

OpenStudy (e.cociuba):

Given two points P(–1, 3) and Q(7, 5) ➊ METHOD ONE Let M name the midpoint between P and Q M has co-ords ( ½(Xp +Xq), ½(Yp + Yq) ) = ( ½(–1 + 7), ½(3 + 5) ) = ( ½(6), ½(8) ) = (3, 4) Let N name the mid point between P and M. N will have co-ords: ( ½(Xp +Xm), ½(Yp + Ym) ) = ( ½(–1 + 3), ½(3 + 4) ) = ( ½(2), ½(7) ) = (1, 3½) ➋ METHOD TWO (which is more efficient ie gets an answer more quickly) Let N name the point that is "one-fourth of the distance from the point P to the point Q along the segment PQ" N will have co-ords: (Xp + ¼(Xq – Xp), Yp + ¼(Yq – Yp) This means that Xn is Xp + one quarter of the difference between Xq and Xp This means that Yn is Yp + one quarter of the difference between Yq and Yp = (–1 + ¼(7 – (–1)), 3 + ¼(5 – 3) ) = (–1 + ¼(8), 3 + ¼(2) ) = (–1 + 2, 3 + ½ ) = (1, 3½) Hope this helps! God Bless You :)

OpenStudy (anonymous):

Awesome! Thank you so much!

OpenStudy (e.cociuba):

Np:)